mikewwe13:

Let ​ f(x) = x^2 + 15x + 56 . What are the zeros of the function? Enter your answers in the boxes. ______ and _________

mikewwe13:

@Vocaloid

dude:

\(x^2 + 15x + 56=0\) We can use the quadratic equation \(\large{\frac{-b\pm \sqrt{b^2-4ac}}{2a}}\)

mikewwe13:

-7 and -8 correct ?

dude:

Right

mikewwe13:

there right ?